# deCIFer: Crystal Structure Prediction from Powder Diffraction Data using Autoregressive Language Models

### v2

@@ Abstract
Novel materials drive progress across applications from energy storage to electronics. Automated characterization of material structures with machine learning methods offers a promising strategy for accelerating this key step in material design. In this work, we introduce an autoregressive language model that performs crystal structure prediction (CSP) from powder diffraction data. The presented model, deCIFer, generates crystal structures in the widely used Crystallographic Information File (CIF) format and can be conditioned on powder X-ray diffraction (PXRD) data. Unlike earlier works that primarily rely on high-level descriptors like composition, deCIFer performs CSP from diffraction data. We train deCIFer on nearly 2.3M unique crystal structures and validate on diverse sets of PXRD patterns for characterizing challenging inorganic crystal systems. Qualitative and quantitative assessments using the residual weighted profile and Wasserstein distance show that deCIFer produces structures that more accurately match the target diffraction data when conditioned, compared to the unconditioned case. Notably, deCIFer can achieve a 94% match rate on unseen data. deCIFer bridges experimental diffraction data with computational CSP, lending itself as a powerful tool for crystal structure characterization and accelerating materials discovery.  [^1] Source code available at: https://github.com/FrederikLizakJohansen/deCIFer
@@
Characterizing the atomic structure in functional materials is fundamental for understanding and optimizing materials for e.g. new energy technologies. Such characterization can be done using X-ray diffraction (XRD), and in particular, powder X-ray diffraction (PXRD) is widely used in materials chemistry and related fields as a main characterization tool Cheetham and Goodwin (2014). PXRD patterns contain information on the arrangement of atoms in a sample. However, direct structure solution (i.e. determining all atomic coordinates) is challenging from PXRD data. Instead, PXRD data analysis is generally done based on a known structural model, while parameters can be refined against experimental data Young (1995). Identifying a model for such analysis is often referred to as fingerprinting, and can be a very challenging task, where chemical intuition as well as extensive database searches are required. Even then, model identification is not always successful, and this hinders further material development. Several advances in machine learning (ML) have shown promise in aiding analysis of diffraction data over the last decade Tatlier (2011); Bunn et al. (2016); Oviedo et al. (2019); Wang et al. (2020), including recent advances in generative models for crystal structure prediction (CSP) Jiao et al. (2023); Mohanty et al. (2024); Antunes et al. (2024). While these conventional ML-aided CSP methods typically explore the structural phase space guided by energy evaluations or high-level descriptors like material composition, emerging data-informed CSP approaches have started to integrate diffraction data directly into the generative process Kjær et al. (2023); Guo et al. (2024); Riesel et al. (2024); Lai et al. (2025). This shift represents a fundamental departure from conventional CSP.
Figure 1: deCIFer performs autoregressive crystal structure prediction (CSP) from PXRD data, optionally guided by tokenized crystal descriptors. PXRD embeddings are prepended to the CIF token sequence, enabling the generation of structurally consistent CIFs directly from diffraction data.
In this work, we introduce deCIFer (Figure 1), a transformer-based model that directly generates Crystallographic Information Files (CIFs) – text files that encode for crystal structures – conditioned on the rich structural information explicitly represented in experimental data, such as PXRD patterns. We thus combine the strength of CSP with the possibility to include experimental data in the workflow. This approach can significantly advance the workflow used for atomic structure characterization based on PXRD methods. Instead of conditioning on the raw scattering data, in this work we learn PXRD embeddings based on a variety of data transformations. These transformations are designed to mimic noise due to experimental variations.
We evaluate the performance of deCIFer on a diverse set of PXRD patterns, demonstrating its robustness to noise and its ability to generate high-quality crystal structures. By using learned conditioning embeddings to extract fine-grained structural features from PXRD patterns, deCIFer is the first of its kind that produces syntactically correct and structurally meaningful CIFs while agreeing with reference PXRD patterns.
@@
    A novel autoregressive transformer model for CSP that incorporates materials science domain knowledge by conditioning on PXRD data to generate viable crystal structures in CIF format. This extends the use of CSP models, as experimental data can be incorporated in the material analysis.
@@
    An efficient conditioning mechanism for autoregressive models when dealing with variable length input.
@@
    High-fidelity simulation of experimental settings for PXRD data with Gaussian noise and peak broadening.
@@
    Systematic evaluation on two large-scale datasets: NOMA [^2] NOMA stands for NOMAD Draxl and Scheffler (2019) OQMD Kirklin et al. (2015) & MJain et al. (2013) Aggregation. & CHILI-100K Friis-Jensen et al. (2024).
Background and related work
PXRD is the most widely used method for structural characterization in materials chemistry. A PXRD pattern contains diffraction peaks, whose position and intensity contain information on the periodic structure in crystalline materials, i.e. their atomic positions and structure symmetry. Most quantitative analysis of PXRD data is done using structure refinement, where the parameters in a structural model is refined against experimental data. This requires a structural starting model of the correct structure type and symmetry. Identifying such a model is often a bottleneck for materials characterization.
@@
The transformer architecture Vaswani et al. (2017) with its capacity to capture long-range dependencies has inspired applications beyond natural language processing, including applications in materials science, where transformer-based large language models (LLMs) have been employed for CSP and property analysis. While LLMs have seen emerging use in automation of chemical syntheses Hocky and White (2022); Szymanski et al. (2023); M. Bran et al. (2024), data extraction Gupta et al. (2022); Dagdelen et al. (2024); Polak and Morgan (2024); Schilling-Wilhelmi et al. (2025), and materials simulation and property prediction Zhang et al. (2024); Rubungo et al. (2024); Jablonka et al. (2024), they are not yet as widely used in tasks like materials design. ChemCrow M. Bran et al. (2024), for instance, is an LLM-powered chemistry search engine designed to automate reasoning tasks in materials design and other domains.
@@
Recent works have used fine-tuning to adapt LLMs for CSP. Gruver et al. (2024) fine-tuned Llama-2 models Touvron et al. (2023) on text-encoded atomistic data, enabling tasks like unconditional generation of stable materials. Similarly, Mohanty et al. (2024) fine-tuned LLaMA-3.1-8B Dubey et al. (2024) using QLoRA Dettmers et al. (2024) for efficient CIF generation conditioned on material composition and space group. In contrast, the work on CrystalLLM Antunes et al. (2024) relies on pre-training alone to generate CIFs. It is trained on an extensive corpus of CIFs representing millions of inorganic compounds.
@@
CrystalLLM relies on composition- and symmetry- level crystal descriptors. Its utility, as most other CSP methods, is thus to generate structures without direct reference to experimental data. deCIFer, on the other hand, conditions on PXRD signals, produces structures agreeing with observed diffraction patterns and address the need for models that bridge predictive power with experimental reality.
@@
In parallel to LLM-based approaches, generative models using diffusion- or flow-based frameworks have also emerged for CSP Jiao et al. (2023); Miller et al. (2024); Zeni et al. (2025). These methods rely on composition or partial structural constraints to guide generation, which requires domain knowledge and can limit direct incorporation of experimental data. While they show promise in producing stable crystal configurations, the gap between purely computational crystal structures and experimental data remains a challenge. The recent diffusion-based framework MatterGen Zeni et al. (2025) addresses some of these challenges by enabling conditioning on a broad range of property constraints, and it can generate structures that are more likely to be synthesizable than previous methods; nevertheless, conditioning on PXRD has not been demonstrated in its current implementation, leaving room for further exploration in integrating experimental data more directly.
@@
Consider a crystal structure in CIF (text input) format that is tokenized into a sequence of length $T_{i}$: ${\mathbf{x}}^{i}=(x^{i}_{1},x^{i}_{2},\dots,x^{i}_{T_{i}})$ (see Appendix A.3 for details on tokenization). The corresponding scattering pattern, which in our case is the PXRD pattern denoted by ${\mathbf{y}}^{i}$, is a continuous-valued vector representing the intensity profile of the scattering pattern. The dataset then consists of pairs of CIFs and PXRD patterns, ${\mathcal{D}}=[({\mathbf{x}}^{i},{\mathbf{y}}^{i})]_{i=1}^{N}$.
Given this dataset, we want to minimize the negative conditional log-likelihood over the training data:
@@
This is achieved using a conditional autoregressive model $f_{\mathbf{\Theta}}(\cdot)$ with trainable parameters ${\mathbf{\Theta}}$ based on the transformer architecture Vaswani et al. (2017). Our autoregressive model, deCIFer, generates structures in CIF format when conditioned with an initial prompt and the PXRD data.
Section 2.1 PXRD Conditioning
The PXRD data can be seen as the structural fingerprint of the structures in CIFs. We use this PXRD data to steer the CSP by using it as a conditioning input.
Following the standard procedure in materials chemistry, for each CIF, we generate the discrete diffraction peak data, given as the set ${\mathcal{P}}=\{(q_{k},i_{k})\}_{k=1}^{n}$ using pymatgen Ong et al. (2013). During the model training, ${\mathcal{P}}$ is transformed into the PXRD data, ${\mathbf{y}}$, using different simulated experimental conditions. Formally, let ${\mathcal{T}}$ be a set of transformations that can be applied to each ${\mathcal{P}}$.
@@
Conditioning model: A multi-layered perceptron (MLP) with trainable parameters ${\mathbf{\Phi}}$ is used to embed the PXRD data into a learnable conditioning vector ${\mathbf{e}}=f_{\mathbf{\Phi}}({\mathbf{y}})\in\mathbb{R}^{D}$. This embedding is inserted at the start of the tokenized CIF sequence as a conditioning signal. Jointly training the conditioning network, $f_{\mathbf{\Phi}}$, and the autoregressive model, $f_{\mathbf{\Theta}}$, enables the integration of structural information from the PXRD profiles in the autoregressive generation of CIFs. This joint training along with transformations to the PXRD data results in the final training objective ${\mathcal{L}}({\mathbf{X}}|{\mathbf{Y}};{\mathbf{\Theta}},{\mathbf{\Phi}})$.
Section 2.2 Sequence packing and Isolating CIFs
@@
To efficiently train on variable-length CIF sequences, we implement a batching strategy during training, inspired by recent sequence-packing methods that prevent cross-contamination and improve throughput Kosec et al. (2021). Our approach concatenates multiple tokenized CIF sequences, each of length $T_{i}$ (excluding the conditioning embedding) into segments of fixed context length, $C$. Formally, consider a sequence $\mathbf{S}=[\mathbf{e}^{1},{\mathbf{t}}^{1}_{1},\dots,{\mathbf{t}}^{1}_{T_{1}},\mathbf{e}^{2},{\mathbf{t}}^{2}_{1},\dots,{\mathbf{t}}^{n}_{k}]$, where $n$ is the last fully or partially included CIF. In this notation ${\mathbf{e}}^{i}$ denotes a $D$-dimensional conditioning embedding for the $i$-th CIF, while ${\mathbf{t}}_{j}^{i}$ is the $D$-dimensional input embedding for $x_{j}^{i}$, the $j$-th token of the $i$-th CIF. We choose $k$ such that $|\mathbf{S}|=C$. If adding another CIF exceeds $C$, that CIF is split at the boundary and continued in the next segment. In practice, we set $C=3076$ based on the available GPU memory to balance throughput and memory constraints, which exceeds the length of the longest tokenized CIFs in the NOMA dataset. Although our method guarantees efficient batch utilization, it does occasionally split exceptionally long CIFs between batches ($\approx\!0.04$% of samples in the NOMA dataset; see Figure 11 in the Appendix). To mitigate this, we shuffle the training set at the start of each epoch so that previously split CIFs are more likely to appear in full in subsequent mini-batches, allowing the model to learn from complete sequences.
@@
To isolate different CIFs in the same sequence, we employ an attention mask ${\mathbf{M}}$, defined such that $M_{kl}=1$ iff tokens $k$ and $l$ belong to the same CIF, and $M_{kl}=0$ otherwise. This results in a block-wise diagonal, upper-triangular, attention matrix as shown in Figure 2. To prevent positional information from leaking across CIF boundaries we also reset positional encodings at the start of each CIF by assigning positions from 00 to $T_{i}-1$ for the $i$-th sequence.
Section 2.3 Evaluation Metrics
We evaluate generated structures using four metrics.
1) Residual weighted profile ($R_{\mathrm{wp}}$) is applied to convolved continuous PXRD profiles, measuring the discrepancy between a reference PXRD profile, ${\mathbf{y}}$, and a generated PXRD profile, ${\mathbf{y}}^{*}$. Formally:
$R_{\mathrm{wp}}=\sqrt{\frac{\sum_{i}w_{i}(y_{i}-y_{i}^{*})^{2}}{\sum_{i}w_{i}y_{i}^{2}}},$ (2)
with all weights $w_{i}=1$ here, following convention.
2) Wasserstein distance (WD) is used to compare the non-convolved, discrete intensity peaks of the PXRD patterns. WD measures the minimal transport cost of moving peaks in the reference set ${\mathcal{P}}=\{(q_{k},i_{k})\}_{k=1}^{n}$ to peaks in the generated set ${\mathcal{P}}^{*}=\{(q_{j}^{*},i_{j}^{*})\}_{j=1}^{m}$. Letting $\pi_{kj}$ be a coupling between intensities at positions $q_{k}$ and $q_{j}^{*}$, we have in 1D:
$\mathrm{WD}({\mathcal{P}},{\mathcal{P}}^{*})=\min_{\pi}\sum_{k=1}^{n}\sum_{j=1}^{m}\pi_{kj}\|q_{k}-q_{j}^{*}\|,$ (3)
where $\sum_{j=1}^{n}\pi_{kj}=i_{k}\text{ and }\sum_{k=1}^{m}\pi_{kj}=i_{j}^{*}$.
3) Match rate (MR) uses StructureMatcher Ong et al. (2013) to assess structural similarity between reference and generated CIFs. Two structures are considered a match if their lattice parameters, atomic coordinates, and space group symmetries are within the defined tolerances. MR is the fraction of matching structures. See Section A.6 in the Appendix for more information.
4) Validity (Val.) ensures the internal consistency of each generated CIF by checking formula consistency, site multiplicity, bond lengths, and space group alignment. A CIF is deemed overall valid only if it passes all four checks. Detailed explanations of these metrics are provided in Section A.5 in the Appendix.
Figure 3 summarizes the evaluation pipeline.
Figure 3: Evaluation pipeline. A reference CIF from the test set is used to generate discrete peaks, $\mathcal{P}=\{(q_{k},i_{k})\}_{k=1}^{n}$. These peaks are then transformed into a PXRD profile $\mathbf{y}=\tau_{\mathrm{fixed}}(\mathcal{P})$, where $\tau_{\mathrm{fixed}}$ is selected according to the dired experimental setting. The resulting PXRD, along with any optional crystal descriptors, is tokenized and passed to deCIFer to produce a new CIF ($\text{CIF}^{*}$). For evaluation, the generated structures are compared using a clean reference transformation $\tau_{0}$ (FWHM$=0.05$, $\sigma_{\mathrm{noise}^{2}}=0$).
@@
Dataset: We use two datasets here. The first, NOMA, is a large-scale compilation of crystal structures curated in CrystaLLM Antunes et al. (2024), which draws from the Materials Project (April 2022) Jain et al. (2013), OQMD (v. 1.5, October 2023) Kirklin et al. (2015), and NOMAD (April 2023) Draxl and Scheffler (2019) databases. The second, CHILI-100K Friis-Jensen et al. (2024), is a large-scale dataset of experimentally determined crystal structures obtained from a curated subset of the Crystallography Open Database (COD) Gražulis et al. (2009). We employ NOMA for both training and testing our models, whereas CHILI-100K is used exclusively for testing. Both datasets are open-source and available for download. [^3] NOMA: github.com/lantunes/CrystaLLM (CC-BY 4.0 licence), CHILI-100K: github.com/UlrikFriisJensen/CHILI (Apache 2.0 licence).
@@
Preprocessing: We begin be applying the standard data preprocessing steps outlined in CrystaLLM before performing additional curation to address complexities such as ionic charges and to ensure consistent structural representation between NOMA and CHILI-100K. The resulting NOMA dataset comprises approximately $2.3$M CIFs, spanning 111110101010 elements, including elements up to atomic number $94$ (excluding polonium, astatine, radon, francium, and radium). Duplicate structures were filtered by selecting the configuration with the lowest volume per formula unit. All structures were converted to a standardized CIF format using the pymatgen library Ong et al. (2013) and all floating point values were rounded to four decimal places. The resulting CHILI-100K dataset consists of $\approx\!8.2$K CIFs, spanning 11118888 elements, including elements up to atomic number $85$. Figures 11 and 12 (in Appendix) show the distribution of elemental compositions, space groups, and more for NOMA and CHILI-100K.
A challenge in NOMA is the disproportionate representation of high-symmetry structures. To account for the skew in the distribution, we derive the space group, a numerical label from $1$ to $230$ that encodes crystal symmetry, from each CIF. We stratify the data by grouping these labels into bins of $10$, e.g., space groups 111110101010, 1111111120202020, etc., and use these strata to split data into $90$% training, $7.5$% validation, and $2.5$% testing subsets while preserving the overall distribution of crystal symmetries.
@@
The dataset was tokenized into a vocabulary of $373$ tokens, encompassing CIF tags, space group symbols, element symbols, numeric digits, punctuation marks, padding tokens and a special conditioning token. For a detailed explanation of the standardization and tokenization processes, refer to Section A.2 and A.3 in the Appendix.
Hyperparameters:

### v3

@@ Abstract
Novel materials drive progress across applications from energy storage to electronics. Automated characterization of material structures with machine learning methods offers a promising strategy for accelerating this key step in material design. In this work, we introduce an autoregressive language model that performs crystal structure prediction (CSP) from powder diffraction data. The presented model, deCIFer, generates crystal structures in the widely used Crystallographic Information File (CIF) format and can be conditioned on powder X-ray diffraction (PXRD) data. Unlike earlier works that primarily rely on high-level descriptors like composition, deCIFer is also able to use diffraction data to perform CSP. We train deCIFer on nearly 2.3M crystal structures and validate on diverse sets of PXRD patterns for characterizing challenging inorganic crystal systems. Qualitative checks and quantitative assessments using the residual weighted profile show that deCIFer produces structures that more accurately match the target diffraction data. Notably, deCIFer can achieve a 94% match rate on test data. deCIFer bridges experimental diffraction data with computational CSP, lending itself as a powerful tool for crystal structure characterization.  [^1] Source code available at: https://github.com/FrederikLizakJohansen/deCIFer
@@
Characterizing the atomic structure in functional materials is fundamental for understanding and optimizing materials for e.g. new energy technologies. Such characterization can be done using X-ray diffraction (XRD), and in particular, powder X-ray diffraction (PXRD) is widely used in materials chemistry and related fields as a main characterization tool (Cheetham and Goodwin, 2014). Several advances in machine learning (ML) have shown promise in aiding analysis of diffraction data over the last decade (Tatlier, 2011; Bunn et al., 2016; Oviedo et al., 2019; Wang et al., 2020), including recent advances in generative models for crystal structure prediction (CSP) (Jiao et al., 2023; Mohanty et al., 2024; Antunes et al., 2024). These conventional ML-aided CSP methods typically explore the structural phase space guided by high-level descriptors like material composition. Emerging experimental data informed CSP approaches have started to integrate diffraction data directly into the generative process (Kjær et al., 2023; Guo et al., 2024; Riesel et al., 2024; Lai et al., 2025). This shift represents a fundamental departure from conventional CSP.
(a) Overview of the deCIFer model.
(b) Example generations using deCIFer.
Figure 1: (a) Overview of the deCIFer model, which performs autoregressive crystal structure prediction (CSP) from PXRD data, optionally guided by tokenized crystal descriptors. PXRD embeddings are prepended to the CIF token sequence, enabling the generation of structurally consistent CIFs directly from diffraction data. (b) Three examples from the NOMA test set showing deCIFer generations, each illustrating a reference structure, the generated structure and their corresponding PXRD profiles.
In this work, we present deCIFer (Figure 1), a transformer-based model that performs CSP using the rich structural information explicitly represented in experimental data, such as PXRD patterns. Our autoregressive language model directly generates the Crystallographic Information Files (CIFs) – text files that encode for crystal structures. This approach can significantly advance the workflow used for atomic structure characterization based on PXRD methods. Furthermore, we mimic the noise in the PXRD as a result of experimental variations using various transformations to the PXRD patterns.
The performance of deCIFer is evaluated on a diverse set of PXRD patterns, demonstrating its robustness to experimental noise and its ability to generate high-quality crystal structures. deCIFer is the first experimental data informed CSP model capable of generating syntactically correct and structurally meaningful CIFs that accurately reproduce reference PXRD patterns.
@@
    Integration of experimental PXRD conditioning into an autoregressive transformer for CSP in CIF format; a capability not demonstrated in existing transformer-based generative models.
@@
    Effective conditioning mechanism for autoregressive models with variable length input.
@@
    Simulation of realistic experimental settings for PXRD with Gaussian noise and peak broadening.
@@
    Comprehensive evaluation on two large-scale datasets: NOMA [^2] NOMA stands for NOMAD (Draxl and Scheffler, 2019) OQMD (Kirklin et al., 2015) & M(Jain et al., 2013b) Aggregation. & CHILI-100K (Friis-Jensen et al., 2024), and comparison to state-of-the-art CSP generative models.
Section 2 Background and Related Work
PXRD: It is arguably the most accessible structural probe in solid-state chemistry. Modern benchtop diffractometers can collect a high-quality scan within minutes and are available in most research and industrial laboratories. A PXRD pattern contains diffraction peaks, whose position and intensity contain information on the periodic structure in crystalline materials, i.e. their atomic positions and structure symmetry. Crucially, the forward simulation of PXRD patterns from CIFs is grounded in well-established scattering theory (West, 2014), enabling realistic modelling.
Most quantitative analysis of PXRD data is done using structure refinement, where the parameters in a structural model is refined against experimental data (Young, 1995). This requires a structural starting model of the correct structure type and symmetry. Identifying a model for such analysis is often referred to as fingerprinting, and can be a very challenging task, where chemical intuition as well as extensive database searches are required. Even then, model identification is not always successful, and this hinders further material development.
@@
CSP with LLMs: Transformer architecture (Vaswani et al., 2017) based large language models (LLMs) have seen emerging use in the automation of chemical syntheses (Hocky and White, 2022; Szymanski et al., 2023; M. Bran et al., 2024), data extraction (Gupta et al., 2022; Dagdelen et al., 2024; Polak and Morgan, 2024; Schilling-Wilhelmi et al., 2025), and materials simulation and property prediction (Zhang et al., 2024; Rubungo et al., 2024; Jablonka et al., 2024). However, they are not yet as widely used in tasks like materials design. ChemCrow (M. Bran et al., 2024), for instance, is an LLM-powered chemistry search engine designed to automate reasoning tasks in materials design and other domains.
@@
Recent works have used fine-tuning to adapt LLMs for CSP. Gruver et al. (2024) fine-tuned Llama-2 models (Touvron et al., 2023) on text-encoded atomistic data, enabling tasks like unconditional generation of stable materials. Similarly, Mohanty et al. (2024) fine-tuned LLaMA-3.1-8B (Dubey et al., 2024) using QLoRA (Dettmers et al., 2024) for efficient CIF generation conditioned on material composition and space group. In contrast, the work on CrystaLLM (Antunes et al., 2024) relies on pre-training alone to generate CIFs. It is trained on an extensive corpus of CIFs representing millions of inorganic compounds.
@@
CrystaLLM relies on composition- and symmetry- level crystal descriptors. Its utility, as most other CSP methods, is thus to generate structures without direct reference to experimental data. deCIFer, on the other hand, conditions on PXRD signals, produces structures agreeing with observed diffraction patterns and address the need for models that bridge predictive power with experimental reality.
@@
CSP with diffusion models: In parallel to LLM-based approaches, generative models using diffusion- or flow-based frameworks have also emerged for CSP (Jiao et al., 2023; Miller et al., 2024; Zeni et al., 2025; Xie et al., 2022). These methods rely on composition or partial structural constraints to guide generation, which requires domain knowledge and can limit direct incorporation of experimental data. While they show promise in producing stable crystal configurations, the gap between purely computational crystal structures and experimental data remains a challenge. The recent diffusion-based framework MatterGen (Zeni et al., 2025) addresses some of these challenges by enabling conditioning on a broad range of property constraints, and it can generate structures that are more likely to be synthesizable than previous methods; nevertheless, conditioning on PXRD has not been demonstrated in its current implementation, leaving room for further exploration in integrating experimental data more directly.
@@
Consider a crystal structure in CIF (text input) format that is tokenized into a sequence of length $T_{i}$: ${\mathbf{x}}^{i}=(x^{i}_{1},x^{i}_{2},\dots,x^{i}_{T_{i}})$ (see Appendix A.3 for details on tokenization). The corresponding scattering pattern, which in our case is the PXRD pattern denoted by ${\mathbf{y}}^{i}$, is a continuous-valued vector representing the intensity profile of the scattering pattern. The dataset then consists of pairs of CIFs and PXRD patterns, ${\mathcal{D}}=[({\mathbf{x}}^{i},{\mathbf{y}}^{i})]_{i=1}^{N}$. Given this dataset, we want to minimize the negative conditional log-likelihood over the training data:
@@
This is achieved using a conditional autoregressive model $f_{\mathbf{\Theta}}(\cdot)$ with trainable parameters ${\mathbf{\Theta}}$ based on the transformer architecture (Vaswani et al., 2017). Our autoregressive model, deCIFer, generates structures in CIF format when conditioned with an initial prompt and the PXRD data.
PXRD Conditioning:
The PXRD data can be seen as the structural fingerprint of the structures in CIFs. We use this PXRD data to steer the CSP by using it as a conditioning input to our model.
Following the standard procedure in materials chemistry, for each CIF, we generate the discrete diffraction peak data, given as the set ${\mathcal{P}}=\{(q_{k},i_{k})\}_{k=1}^{n}$ using pymatgen (Ong et al., 2013). During the model training, ${\mathcal{P}}$ is transformed into the PXRD data, ${\mathbf{y}}$, using different simulated experimental conditions. Formally, let ${\mathcal{T}}$ be a set of transformations that can be applied to each ${\mathcal{P}}$.
@@
A multi-layered perceptron (MLP) $f_{\mathbf{\Phi}}({\mathbf{y}})$ with trainable parameters ${\mathbf{\Phi}}$ is used to embed the PXRD data into a learnt vector ${\mathbf{e}}=f_{\mathbf{\Phi}}({\mathbf{y}})\in\mathbb{R}^{D}$. This embedding is inserted at the start of the tokenized CIF sequence as a conditioning signal. Joint training of the conditioning network, $f_{\mathbf{\Phi}}$, and the autoregressive model, $f_{\mathbf{\Theta}}$, enables the integration of structural information from the PXRD profiles during the generation of CIFs. This joint training along with transformations to the PXRD data results in the final training objective ${\mathcal{L}}({\mathbf{X}}|{\mathbf{Y}};{\mathbf{\Theta}},{\mathbf{\Phi}})$.
Sequence Packing and Isolating CIFs:
@@
To effectively train on variable-length CIF sequences, we implement a batching strategy during training, inspired by recent sequence-packing methods that prevent cross-contamination and improve throughput (Kosec et al., 2021). Our approach concatenates multiple tokenized CIF sequences, each of length $T_{i}$ (excluding the conditioning embedding) into segments of fixed context length, $C$. Formally, consider a sequence $\mathbf{S}=[\mathbf{e}^{1},{\mathbf{t}}^{1}_{1},\dots,{\mathbf{t}}^{1}_{T_{1}},\mathbf{e}^{2},{\mathbf{t}}^{2}_{1},\dots,{\mathbf{t}}^{n}_{k}]$, where $n$ is the last fully or partially included CIF. In this notation ${\mathbf{e}}^{i}$ denotes a $D$-dimensional conditioning embedding for the $i$-th CIF, while ${\mathbf{t}}_{j}^{i}$ is the $D$-dimensional input embedding for $x_{j}^{i}$, the $j$-th token of the $i$-th CIF. We choose $k$ such that $|\mathbf{S}|=C$. If adding another CIF exceeds $C$, that CIF is split at the boundary and continued in the next segment. In practice, we set $C=3076$ based on the available GPU memory to balance throughput and memory constraints, which exceeds the length of the longest tokenized CIFs in the NOMA dataset. Although our method guarantees efficient batch utilization, it does occasionally split exceptionally long CIFs between batches ($\approx\!0.04$% of samples in the NOMA dataset; see Figure 10 in the Appendix). To mitigate this, we shuffle the training set at the start of each epoch so that previously split CIFs are more likely to appear in full in subsequent mini-batches, allowing the model to learn from complete sequences.
@@
To isolate different CIFs in the same sequence, we employ an attention mask ${\mathbf{M}}$, defined such that $M_{kl}=1$ iff tokens $k$ and $l$ belong to the same CIF, and $M_{kl}=0$ otherwise. This results in a block-wise diagonal, upper-triangular, attention matrix as shown in Figure 8 (Appendix). To prevent positional information from leaking across CIF boundaries we also reset positional encodings at the start of each CIF by assigning positions from 00 to $T_{i}-1$ for the $i$-th sequence.
@@
We use two large-scale public datasets here. The first, NOMA, is a large-scale compilation of crystal structures curated in CrystaLLM (Antunes et al., 2024), which draws from the Materials Project (April 2022) (Jain et al., 2013b), OQMD (v. 1.5, October 2023) (Kirklin et al., 2015), and NOMAD (April 2023) (Draxl and Scheffler, 2019) databases. The second, CHILI-100K (Friis-Jensen et al., 2024), is a large-scale dataset of experimentally determined crystal structures obtained from a curated subset of the Crystallography Open Database (COD) (Gražulis et al., 2009). We employ NOMA for both training and testing our models, whereas CHILI-100K is used exclusively for testing. Both datasets are open-source and available for download. [^3] NOMA: github.com/lantunes/CrystaLLM (CC-BY 4.0 licence), CHILI-100K: github.com/UlrikFriisJensen/CHILI (Apache 2.0 licence).
Preprocessing:
@@
We perform standard data preprocessing steps outlined in CrystaLLM before performing additional curation to address complexities such as ionic charges and to ensure consistent structural representation between NOMA and CHILI-100K. The resulting NOMA dataset comprises approximately $2.3$M CIFs, spanning 111110101010 elements, including elements up to atomic number $94$ (excluding polonium, astatine, radon, francium, and radium). Duplicate structures were filtered by selecting the configuration with the lowest volume per formula unit. All structures were converted to a standardized CIF format using the pymatgen library (Ong et al., 2013) and all floating point values were rounded to four decimal places. The resulting CHILI-100K dataset consists of $\approx\!8.2$K CIFs, spanning 11118888 elements, including elements up to atomic number $85$. Figures 10 and 11 (in Appendix) show the distribution of elemental compositions, space groups, and more attributes for the two datasets.
Additionaly, a known challenge in datasets like NOMA is the disproportionate representation of high-symmetry structures, which can skew model learning and evaluation (Davariashtiyani et al., 2024; Zhang et al., 2023). To mitigate this, we stratify the dataset during the train/validation/test split based on space group labels. For details on stratification, see Section A.9.
@@
Finally, CIFs from both datasets were tokenized into a vocabulary of $373$ tokens, encompassing CIF tags, space group symbols, element symbols, numeric digits, punctuation marks, padding tokens and a special conditioning token. For a detailed explanation of the standardization and tokenization processes, refer to Section A.2 and A.3 in the Appendix.
Model Hyperparameters: